OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0

OpenStudy (shamil98):

3x + 5y =3 x + 2y = 0 x = -2y 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3 x + 2(-3) = 0 x -6 = 0 x = 6

OpenStudy (anonymous):

Thanks :)